OpenStudy (anonymous):

how do you show that the maximum speed of an oscillation in simple harmonic motion is the product of the angular velocity and the area?

OpenStudy (anonymous):

It is not the angular velocity times the area. It is the angular velocity times the amplitude. The equation for simple harmonic motion is \[x(t) = A \cos(\omega t) _______ (1)\] A is the amplitude, the maximum value of x omega is the angular or circular frequency From this you can get the velocity by differentiating with respect to t: \[v(t) = dx/dt = - \omega A \sin(\omega t) ________ (2)\] From (2) you can see that the amplitude of the velovity, which is the maximim value of the speed is omega * A